### v1

# Hilbert property for low-genus families and degree-one del Pezzo surfaces

Abstract.
We prove that the Hilbert property is satisfied by certain del Pezzo surfaces of degree one and Picard rank 1 over fields finitely generated over $\mathbb{Q}$. We generalize results of the first author on elliptic surfaces and employ constructions used by Desjardins and the third author to prove density of rational points. Our results are the first on the Hilbert property for minimal del Pezzo surfaces of degree one without a conic fibration.
2020 Mathematics Subject Classification:
14G05, 14J20 (14G25, 14G27, 11G35).
1. Introduction
This work is primarily concerned with the following long-standing question of Colliot-Thélène and Sansuc [4] about the abundance of rational points on unirational varieties.
Question (Colliot-Thélène–Sansuc).
Does every projective unirational variety over a Hilbertian field satisfy the Hilbert property?
For the definitions of Hilbertian fields and the Hilbert property, see Section 2.1.
The Hilbert property takes its name from Hilbert’s Irreducibility Theorem on specializations of polynomials, which establishes the Hilbert property for $\mathbb{A}^{n}_{K}$ for $K$ a number field. The Hilbert property is closely connected to the inverse Galois problem: if all unirational varieties over a number field $K$ satisfy the Hilbert property, then for any finite group $G$, there exists a Galois extension $L/K$ with $\operatorname{Gal}(L/K)\cong G$ [24, Proof of Thm. 3.5.9].
In this paper we generalize a result of the first author in [7, Thm. 1.2] proving the Hilbert property for surfaces over a number field with multiple genus-$1$ fibrations which are sufficiently transversal (Definition 2.5). In our adaptation of this theorem, we allow one of the elliptic fibrations to be replaced by a low-genus family (Definition 2.7).
Theorem 1.1.
Let $S$ be a smooth projective geometrically connected surface over a field $K$ which is finitely generated over $\mathbb{Q}$. Suppose given genus-$1$ fibrations $\pi_{i}:S\rightarrow\mathbb{P}^{1}$, $1\leq i\leq n$ and a low-genus family $\phi:X\rightarrow S$ with geometrically integral generic fiber. Let $F\subset S$ denote the union of divisors on $S$ which are vertical with respect to each $\pi_{i}$ and do not intersect a general elements of $\phi$ transversally. If $S\setminus F$ is algebraically simply connected and $X(K)\subset X$ is Zariski-dense, then $S$ satisfies the Hilbert property over $K$.
Recall that a smooth variety $V$ is algebraically simply connected if any finite surjective morphism $V^{\prime}\rightarrow V$ of degree at least $2$ is ramified.
We apply Theorem 1.1 to certain families of del Pezzo surfaces (smooth projective varieties $X$ of dimension $2$ with ample anticanonical class $K_{X}$). It may be that all del Pezzo surfaces with a rational point are unirational and so within the remit of the above question [21, Rem. 9.4.11]. Since del Pezzo surfaces of degree one always possess a rational point (the base-point of the anticanonical linear system), we are thus motivated to prove that they satisfy the Hilbert property.
The arithmetic complexity of del Pezzo surfaces is notionally governed by the degree $d=(-K_{X})^{2}\in\{1,\dots,9\}$. Let $X/k$ be a degree-$d$ del Pezzo surface over a Hilbertian field $k$ with a rational point. For $d\geq 5$, such $X$ is rational [21, Thm. 9.4.29]; as the Hilbert property is a birational invariant of smooth varieties, Hilbert’s original result then implies the Hilbert property. The Hilbert property is established for $d=4$ in work of the second author [25]. Previous joint work of the authors [9] establishes the Hilbert property when $d=2$ and $X/k$ contains a rational point outside a closed subset of $X$. This result immediately implies the Hilbert property for $d=3$ [9, Cor. 3.7].
When $d=1$, the first two authors [8] proved the Hilbert property when $k$ is a number field and $X$ possesses a general conic fibration. Little is known about rational points on minimal degree-one del Pezzo surfaces with no conic fibration, which are those of Picard rank 1. Desjardins and the third author [11] proved density of rational points for an infinite family of del Pezzo surfaces of degree one. This was generalized by Nijgh [20]. Salgado and van Luijk [22] proved that the collection of $\mathbb{Q}$-rational del Pezzo surfaces with dense rational points is dense in a parameter space for real del Pezzo surfaces, following earlier work over $\mathbb{Q}$ of Várilly-Alvarado [29], Ulas [26, 27], Ulas and Togbé [28] and Jabara [15].
The primary result of this paper is the following.
Theorem 1.2.
Let $K$ be a field finitely generated over $\mathbb{Q}$. Let $a_{4}(u)=au+b,a_{6}(u)=cu^{2}+du+e\in K[u]$ be polynomials and $f(t):=f_{3}t^{3}+f_{2}t^{2}+f_{1}t+f_{0}\in K[t]$ be a cubic polynomial. Let $S\subset\mathbb{P}(2,3,1,1)$ be the surface given by
$S:y^{2}=x^{3}+a_{4}(f(z/w))xw^{4}+a_{6}(f(z/w))w^{6}.$
Assume that $S$ is smooth, hence a del Pezzo surface of degree one. Let $\rho:\mathcal{E}=\text{Bl}_{\mathcal{O}}S\rightarrow S$ be the blowup of $S$ at the base-point $\mathcal{O}=[1:1:0:0]$ of $|{-K_{S}}|$ and $\pi:\mathcal{E}\rightarrow\mathbb{P}^{1}$ be the elliptic fibration given by projection to $[z:w]$. Suppose that there is $P=[x_{0}:y_{0}:z_{0}:w_{0}]\in S(K)$ with $w_{0}\neq 0$, $3z_{0}+2f_{2}w_{0}/f_{3}\neq 0$ and $f(t)-f(z_{0}/w_{0})$ separable such that $\rho^{-1}(P)$ is non-torsion on its fiber of $\pi$. Then $S$ satisfies the Hilbert property over $K$.
Remark 1.3.
By [20, Thm. 3.1], the existence of a point $P$ as in Theorem 1.2 is equivalent to Zariski-density of $S(K)$ since $K$ is finitely generated over $\mathbb{Q}$.
Remark 1.4.
Every del Pezzo surface of degree 1 can be written as a smooth weighted sextic in $\mathbb{P}(2,3,1,1)$, which, over a field $k$ with $\operatorname{char}(k)\neq 2,3$, is of the form
$y^{2}=x^{3}+a_{4}x+a_{6},$
where $a_{i}\in k[z,w]$ is homogeneous of degree $i$. Conversely, each such surface is a del Pezzo surface of degree 1. The family in Theorem 1.2 is thus characterized by $f$.
The proof of Theorem 1.2 makes use of a geometric construction, present in the aforementioned works of Desjardins and the third author [11] and Nijgh [20], which produces a low-genus family on $\mathcal{E}$ sufficiently transversal to $\pi$ to prove density of $S(K)$.
Lastly, we show that, as $a_{4}$ and $a_{6}$ and $f$ vary, the collection of $S$ as in Theorem 1.2 with non-thin rational points and Picard rank $1$ is itself non-thin.
Theorem 1.5.
For any field $K$ finitely generated over $\mathbb{Q}$, there is a non-thin set of parameters $(a,b,c,d,e,f_{0},\ldots,f_{3})\in K^{9}$ such that $S$ as in Theorem 1.2 is of Picard rank $1$ and satisfies the Hilbert property.
1.1. Acknowledgements
The authors thank Daniel Loughran for useful discussions and both the University of Bath and UniDistance Suisse for their hospitality during two week-long workshops. We also thank David Loeffler for financial support during the workshop at UniDistance Suisse. This work was supported by a Focused Research Grant from the Heilbronn Institute for Mathematical Research. The second author is supported by the University of Bristol and the Heilbronn Institute for Mathematical Research. Over the course of this project, the third author was supported by UKRI Fellowship MR/T041609/2, ERC Horizon 2020 Consolidator Grant ID 101001051, and MSCA Postdoctoral Fellowship 101148712.
[Uncaptioned image] Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
2. Low-genus families and the Hilbert property
In this section we prove Theorem 1.1 on the Hilbert property for certain elliptic families. We begin with an overview of the Hilbert property.
2.1. Hilbert property
Definition 2.1.
Let $X$ be a variety over a field $k$. We say that $A\subset X(k)$ is thin if there exist finitely many generically finite dominant morphisms $f_{i}:Y_{i}\rightarrow X$, $i=1,\dots,r$, each of degree at least $2$, such that $X(k)\setminus\bigcup_{i=1}f_{i}(Y_{i}(k))$ is not Zariski-dense in $X$.
Without loss of generality, we can (and do) assume that each $Y_{i}$ is normal and geometrically integral, and each $f_{i}$ is finite [25, Rems. 2.2, 2.3].
Definition 2.2.
We say that a variety $X$ over a field $k$ satisfies/has the Hilbert property (over $k$) if $X(k)$ is not thin. We say that a field $k$ is Hilbertian if there exists a $k$-variety $X$ with the Hilbert property (equivalently, if $\mathbb{P}^{1}_{k}$ has the Hilbert property).
Example 2.3.
Number fields are Hilbertian by Hilbert’s Irreducibility Theorem; more generally and crucially for us, any field finitely generated over $\mathbb{Q}$ is Hilbertian [12, §13.3]. Local fields, finite fields and algebraically closed fields are not Hilbertian.
We now move on to low-genus families.
2.2. Low-genus families
Definition 2.4.
Let $\pi:Y\rightarrow B$ be a surjective flat morphism of varieties over a field $k$.
  1. (1)
    For $g\in\mathbb{Z}_{\geq 0}$, we say that $\pi$ is a genus-$g$ fibration if the generic fiber of $\pi$ is a smooth curve of genus $g$.
  2. (2)
    We say that $\pi$ is a low-genus fibration if it is a genus-$g$ fibration for $g\in\{0,1\}$.
  3. (3)
    We say that $\pi$ is an elliptic fibration if it is a genus-$1$ fibration with a section. If further $Y$ has dimension $2$, then we call $Y$ an elliptic surface over $B$.
Before defining low-genus families, we first define transversal intersection and vertical/horizontal ramification of low-genus fibrations.
Definition 2.5.
Let $X$ be a variety over a field $k$ and let $C$ and $D$ be two divisors on $X$. We say that $C$ and $D$ intersect transversally if the pullback of $C\cap D$ to the desingularization of $C$ contains a reduced point.
Definition 2.6.
Let $\pi\colon Y\rightarrow B$ be a low-genus fibration over a curve $B$ and $\psi:Z\rightarrow Y$ be a cover with branch divisor $D$. We say that $\psi$ is vertically ramified with respect to $\pi$ if $D$ is contained in a union of fibers of $\pi$ (i.e. $D$ is vertical with respect to $\pi$) and horizontally ramified otherwise.
We are now ready to define low-genus families.
Definition 2.7.
Let $X$ be a variety over a field $k$. A low-genus family on $X$ is a surjective morphism $\phi:Y\rightarrow X$ from a variety $Y$ with a low-genus fibration $\pi:Y\rightarrow B$. The elements of the family are the images $\phi(F)$ for $F$ a fiber of $\pi$.
The following result reduces the proof of Theorem 1.1 to a special case.
Theorem 2.8.
Let $S$ be a surface over a field $K$ finitely generated over $\mathbb{Q}$, and let $\pi_{1},\ldots,\pi_{n}:S\to\mathbb{P}^{1}$ be $n\geq 1$ distinct genus-$1$ fibrations. Denote by $\mathcal{C}$ the set of finite covers $\psi:Y\rightarrow S$ of degree at least $2$ which are vertically ramified with respect to each $\pi_{i}$. If, for every finite subset $\{\psi_{j}:Y_{j}\to S,i=j,\ldots,r\}\subset\mathcal{C}$, the set
$S(K)\setminus\cup_{j=1}^{r}\psi_{j}(Y_{j}(K))$
is Zariski-dense in $S$, then $S$ satisfies the Hilbert property.
Proof.
The proof of [7, Thm. 1.1] shows that, for $K$ a number field, given covers $\psi_{j}:Z_{j}\rightarrow S$, $j=1,\dots,s$ of degrees $\geq 2$ such that $S(K)\setminus\bigcup_{j=1}^{s}\psi_{j}(Z_{j}(K))$ is not dense, there is a subset $\{j_{1},\ldots,j_{t}\}\subset\{1,\ldots,s\}$ such that $\psi_{j_{i}}:Z_{j_{i}}\rightarrow S$ is contained in $\mathcal{C}$ for all $j_{i}\in\{j_{1},\ldots,j_{t}\}$ and $S(K)\setminus\bigcup_{i=1}^{t}\psi_{j_{i}}(Z_{j_{i}}(K))$ is not dense (the simply connected hypothesis in loc. cit. ensures that $\mathcal{C}=\emptyset$ in that setting). Since this non-density would contradict our hypothesis, we deduce the result for number fields. To deduce the same for $K$ finitely generated over $\mathbb{Q}$, note that the Mordell–Weil theorem, Faltings’ theorem and Merel’s theorem, which are the key ingredients in the proof of loc. cit. in the setting of number fields, all admit generalizations to such $K$: the generalization of Mordell–Weil follows from the Lang–Néron theorem (see [5, Cor. 7.2]), the generalization of Faltings’ theorem follows from the work of McQuillan [19], and the generalization of Merel’s theorem is established by Colliot-Thélène in [11, Appendix]. ∎
We are now ready to prove Theorem 1.1.
Proof of Theorem 1.1.
Let $\mathcal{C}$ be the set in Theorem 2.8, and let $\{\psi_{i}:V_{i}\rightarrow S,i=1,\dots,s\}$ be a finite subset of $\mathcal{C}$. We will show that $S(K)\setminus\bigcup_{i=1}^{r}\psi_{i}(V_{i}(K))$ is Zariski-dense in $S$, so that we can deduce the result from Theorem 2.8. For each $i\in\{1,\dots,s\}$, denote by $F_{i}$ the fiber product $X\times_{\phi,\psi_{i}}V_{i}$. Let $X_{\eta}$ (resp. $F_{i,\eta}$) be the generic fiber of the low-genus fibration $X\rightarrow B$ (resp. $F_{i}\to X\rightarrow B$). We first show that for all $i\in\{1,\dots,s\}$, the curve $F_{i,\eta}$ is integral over $K(B)$. We have the following commutative diagram.
[…]
Since $S$ is a smooth surface and $V_{i}$ is normal, the morphism $\psi_{i}$ is flat by miracle flatness [1, Tag 00R4]. It follows that $F_{i,\eta}\to X_{\eta}$ is finite flat as well (both finiteness and flatness being preserved under base change), with generic fiber $\operatorname{Spec}K(X)\times_{\operatorname{Spec}K(S)}\operatorname{Spec}K(V_{i})\cong\operatorname{Spec}(K(X)\otimes_{K(S)}K(V_{i}))$. Since $\phi$ has geometrically integral generic fiber, $K(X)\otimes_{K(S)}K(V_{i})$ is a field: the extensions $K(X)/K(S)$ and $K(V_{i})/K(S)$ are regular and algebraic respectively, hence linearly disjoint. Thus the flat morphism $F_{i,\eta}\to X_{\eta}$ has integral generic fiber, and thus $F_{i,\eta}$ is also integral [17, Prop. 4.3.8].
For each $i\in\{1,\ldots,s\}$, denote by $D_{i}\subset S$ the branch locus of $\psi_{i}$, which is a divisor by Nagata–Zariski purity [1, Tag 0BMB]. We claim that a general element of the low-genus family intersects each $D_{i}$ transversally: otherwise, there is at least one $i_{0}$ such that $D_{i_{0}}$ belongs to $F$, but then the restriction of $\psi_{i_{0}}$ over $S\setminus F$ is an unramified cover of degree $\geq 2$, contradicting that $S\setminus F$ is algebraically simply connected. Then there is at least one smooth point on $F_{i,\eta}$ which is ramified over $X_{\eta}$ (cf. [6, p. 595]), hence $F_{i,\eta}$ is an integral $K(B)$-curve ramified over $X_{\eta}$. We may spread out over some open $U\subset B$ to a morphism $\xi_{i}\colon F_{i,U}\to X_{U}$ such that, for all $P\in U$, the curve $F_{i,P}$ is a ramified cover of $X_{P}$.
Since $K$ is finitely generated over $\mathbb{Q}$ and $X(K)$ is Zariski-dense, there is a non-empty Zariski open $V\subset X_{U}$ such that, for each $v\in V(K)$, the fiber $\pi^{-1}(\pi(v))$ has infinitely many $K$-rational points; see [11, Appendix] for the genus-$1$ case (the genus-0 case is elementary, as any smooth curve of genus 0 with one rational point has infinitely many). By assumption, $V(K)$ is Zariski-dense in $V$, and we deduce Zariski-density in $U$ of
$\mathcal{A}:=\{u\in U(K):\#\pi^{-1}(u)(K)=\infty\}.$
When $g=1$, it follows from Lang–Néron that, for each $u\in\mathcal{A}$, only finitely many points in $\pi^{-1}(u)(K)$ can lift to $F_{i,U}$ for some $i$, hence $\pi^{-1}(u)(K)\setminus\cup_{i=1}^{s}\xi_{i}(F_{i,U}(K))$ is dense in $\pi^{-1}(u)$; the same conclusion holds for $g=0$ since $\pi^{-1}(u)$ has the Hilbert property, hence
$\mathcal{A}^{\prime}:=\bigcup_{u\in\mathcal{A}}\pi^{-1}(u)(K)\setminus\cup_{i=1}^{s}\xi_{i}(F_{i,U}(K))$
is Zariski-dense in $X_{U}$ in either case. Thus $\phi(\mathcal{A}^{\prime})$ is Zariski-dense in $S$. Since $\phi(\mathcal{A}^{\prime})$ is contained in $S(K)\setminus\cup_{i=1}^{s}\psi_{i}(V_{i}(K))$, this concludes the proof. ∎
3. Low-genus families on degree-one del Pezzo surfaces
We now introduce the family of degree-one del Pezzo surfaces studied by Nijgh [20]. Throughout this section, let $K$ be a field that is finitely generated over $\mathbb{Q}$.
Definition 3.1.
Let $f(t)\in K[t]$ be a cubic polynomial and define
$a_{4}(t):=at+b,\quad a_{6}(t):=ct^{2}+dt+e\in K[t].$
Let $S\subset\mathbb{P}(2,3,1,1)$ be the surface given by
$S:y^{2}=x^{3}+a_{4}(f(z/w))xw^{4}+a_{6}(f(z/w))w^{6}.$
The surface $S$ is smooth if and only if the sextic branch curve $x^{3}+a_{4}(f(z/w))xw^{4}+a_{6}(f(z/w))w^{6}=0$ in $\mathbb{P}(2,1,1)$ is. We assume that $S$ is smooth from now on, hence a del Pezzo surface of degree one. Recall that we have $\rho:\mathcal{E}\rightarrow S$ the blowup of $S$ at the base-point $\mathcal{O}=[1:1:0:0]$ of $|{-K_{S}}|$ and $\pi:\mathcal{E}\rightarrow\mathbb{P}^{1}$ the associated elliptic fibration.
Note 3.2.
The above family corresponds to the one studied by Nijgh in [20] via linear change of variables. Further, it strictly contains the family
$y^{2}=x^{3}+cz^{6}+dz^{3}w^{3}+ew^{6},$
studied in [11], recovered by the choices $f(t)=t^{3}$, $a_{4}=0$ and $a_{6}=ct^{2}+dt+e$.
Consider the regular map $\theta:S\to\mathbb{P}^{3},[x:y:z:w]\mapsto[xw:y:w^{3}f(z/w):w^{3}]$. The image $W:=\theta(S)$ is a singular cubic surface in $\mathbb{P}^{3}$ with equation
$W:X_{1}^{2}X_{3}=X_{0}^{3}+a_{4}(X_{2}/X_{3})X_{0}X_{3}^{2}+a_{6}(X_{2}/X_{3})X_{3}^{3}.$
Lemma 3.3.
The following hold.
  1. (1)
    The singular locus of $W$ is $\{[0:\pm\sqrt{c}:1:0]\}$.
  2. (2)
    The types of singularities on $W$ are as follows:
    1. (a)
      If $c\neq 0$, we obtain two $A_{2}$ singularities.
    2. (b)
      If $c=0$ and $a\neq 0$, we obtain one $A_{5}$ singularity.
    3. (c)
      If $c=a=0$, we obtain one $E_{6}$ singularity.
  3. (3)
    The cubic surface $W$ is not a cone.
Proof.
  1. (1)
    Let $W^{\text{aff}}\subset\mathbb{A}^{3}$ and $S^{\text{aff}}\subset\mathbb{A}^{3}$ be the restrictions of $W$ and $S$ to the affine charts $X_{3}\neq 0$ and $w\neq 0$, respectively given by $x_{1}^{2}=x_{0}^{3}+a_{4}(x_{2})x_{0}+a_{6}(x_{2})$ and $y^{2}=x^{3}+a_{4}(f(z))x+a_{6}(f(z))$. Note that $S^{\text{aff}}=\phi^{-1}(W^{\text{aff}})$ for the flat map
    $\phi:\mathbb{A}^{3}\to\mathbb{A}^{3},\quad(x,y,z)\mapsto(x,y,f(z)),$
    and the induced map $S^{\text{aff}}\to W^{\text{aff}}$ is also flat by preservation of flatness under base-change. Now, any variety $V$ with a flat surjective morphism $V^{\prime}\to V$ from a smooth variety $V^{\prime}$ is smooth (see [18, Thm. 23.7]), hence $W^{\text{aff}}$ is smooth. Then all singularities of $W$ lie on $X_{3}=0$.
    When $X_{3}=0$, the equation for $W$ also gives $X_{0}=0$. The point $[0:1:0:0]$ is smooth, as $(\partial F_{W}/\partial X_{3})(0,1,0,0)=-1$, with $F_{W}:=X_{0}^{3}+a_{4}(X_{2}/X_{3})X_{0}X_{3}^{2}+a_{6}(X_{2}/X_{3})X_{3}^{3}-X_{1}^{2}X_{3}$. The remaining points on the line $X_{0}=X_{3}=0$ are covered by the affine chart $X_{2}\neq 0$, where the equation becomes $x_{1}^{2}x_{3}=x_{0}^{3}+a_{4}^{\prime}(x_{3})x_{0}x_{3}+a_{6}^{\prime}(x_{3})x_{3},$ with $a_{4}^{\prime},a_{6}^{\prime}\in k[t]$ defined by $a_{4}^{\prime}(t)=ta_{4}(1/t)$ and $a_{6}^{\prime}(t)=t^{2}a_{6}(1/t)$. We rewrite the equation as $x_{1}^{2}x_{3}-cx_{3}=O(x_{0},x_{3})^{2}$ and deduce that the only singular points with $x_{0}=x_{3}=0$ are $(0,\pm\sqrt{c},0)$.
  2. (2)
    To determine the singularity type/types, we follow [2, §2].
    1. (a)
      First assume that $c\neq 0$.
      1. (i)
        If further $a\neq 0$, then
        $F_{W}\mathopen{}\mathclose{{\left(\frac{2\sqrt{c}}{a}\mathopen{}\mathclose{{\left(X_{0}-\frac{d}{2\sqrt{c}}-X_{2}}}\right),\frac{1}{2}(-X_{2}+X_{3}),\frac{1}{2\sqrt{c}}(X_{2}+X_{3}),X_{1}}}\right)$
        $=X_{0}X_{1}X_{3}+G_{W}(X_{0},X_{1},X_{2}),$
        $G_{W}(X_{0},X_{1},X_{2})$ $=\frac{8c\sqrt{c}}{a^{3}}X_{0}^{3}-\frac{12cd}{a^{3}}X_{0}^{2}X_{1}+\frac{2a^{2}b\sqrt{c}+6\sqrt{c}d^{2}}{a^{3}}X_{0}X_{1}^{2}$
        $+\frac{a^{3}e-a^{2}bd-d^{3}}{a^{3}}X_{1}^{3}-\frac{24c\sqrt{c}}{a^{3}}X_{0}^{2}X_{2}$
        $+\frac{a^{3}+24cd}{a^{3}}X_{0}X_{1}X_{2}+\frac{-2a^{2}b\sqrt{c}-6\sqrt{c}d^{2}}{a^{3}}X_{1}^{2}X_{2}$
        $+\frac{24c\sqrt{c}}{a^{3}}X_{0}X_{2}^{2}+\frac{-a^{3}-12cd}{a^{3}}X_{1}X_{2}^{2}-\frac{8c\sqrt{c}}{a^{3}}X_{2}^{3}.$
        Since $G_{W}(0,0,1)\neq 0$, one of the singularities is $A_{2}$ by [2, Lem. 3], hence so is the other singularity by symmetry.
      2. (ii)
        If instead $a=0$, then
        $F_{W}\mathopen{}\mathclose{{\left(X_{2},\frac{1}{2}\mathopen{}\mathclose{{\left(-X_{0}+\frac{d}{2\sqrt{c}}X_{1}+X_{3}}}\right),\frac{1}{2\sqrt{c}}\mathopen{}\mathclose{{\left(X_{0}-\frac{d}{2\sqrt{c}}X_{1}+X_{3}}}\right),X_{1}}}\right)$
        $=X_{0}X_{1}X_{3}+G_{W}(X_{0},X_{1},X_{2}),$
        $G_{W}(X_{0},X_{1},X_{2})$ $=\frac{d}{2\sqrt{c}}X_{0}X_{1}^{2}+\frac{4ce-d^{2}}{4c}X_{1}^{3}+bX_{1}^{2}X_{2}+X_{2}^{3}.$
        Since $G_{W}(0,0,1)\neq 0$, we reach the same conclusion as in (i).
    2. (b)
      Next suppose that $c=0$ and $a\neq 0$. Then
      $F\mathopen{}\mathclose{{\left(\frac{1}{a}(X_{0}-dX_{1}),X_{2},X_{3},X_{1}}}\right)=X_{0}X_{1}X_{3}+G_{W}(X_{0},X_{1},X_{2},X_{3}),$
      $G_{W}(X_{0},X_{1},X_{2})$ $=\frac{1}{a^{3}}X_{0}^{3}-\frac{3d}{a^{3}}X_{0}^{2}X_{1}+\frac{a^{2}b+3d^{2}}{a^{3}}X_{0}X_{1}^{2}$
      $+\frac{a^{3}e-a^{2}bd-d^{3}}{a^{3}}X_{1}^{3}-X_{1}X_{2}^{2}.$
      Since $g_{1}(X_{1}):=G_{W}(0,X_{1},1)$ has a zero of order $1$ at $X_{1}=0$ and $g_{0}(X_{0}):=G_{W}(X_{0},0,1)$ has a zero of order $3$ at $X_{0}=0$, the singularity is $A_{5}$ [2, Lem. 3].
    3. (c)
      Now assume that $c=a=0$; by smoothness of $S$, we find $d\neq 0$. Then
      $F\mathopen{}\mathclose{{\left(X_{2},X_{1},X_{3},\frac{1}{\sqrt{d}}X_{0}}}\right)=X_{3}X_{0}^{2}+G_{W}(X_{0},X_{1},X_{2}),$
      $G_{W}(X_{0},X_{1},X_{2}):=X_{2}^{3}+\frac{1}{d\sqrt{d}}X_{0}^{2}\mathopen{}\mathclose{{\left(eX_{0}+b\sqrt{d}X_{2}}}\right)-\frac{1}{\sqrt{d}}X_{0}X_{1}^{2}.$
      Since $G_{W}(0,X_{1},X_{2})=X_{2}^{3}$, the singularity is $E_{6}$ by [2, Lem. 4].
  3. (3)
    If $W$ were a cone, then since all of its singularities are isolated by the above, it would have to be a cone over a smooth plane cubic $C$. Denote by $\phi:W\dashrightarrow C$ the natural projection. The composition $\phi\circ\theta:S\dashrightarrow C$ is a dominant rational map from $S$ to the curve $C$ of genus $1$. However, since $S$ is a geometrically rational surface, such a map cannot exist and this is a contradiction. ∎
Let $W^{o}:=W\setminus\{[0:\pm\sqrt{c}:1:0]\}$ be the smooth locus of $W$, and set $S^{o}:=\theta^{-1}(W^{o})$.
Definition 3.4.
We define the following curves:
  1. (1)
    For $P\in S^{o}$, we define $C_{P}:=\theta^{-1}(T_{\theta(P)}W\cap W)$ for $T_{\theta(P)}W$ the tangent plane.
  2. (2)
    For $P\in\mathcal{E}$, we set $E_{P}:=\pi^{-1}(\pi(P))\subset\mathcal{E}$.
  3. (3)
    For $P\in S\setminus\{\mathcal{O}\}$, we define $D_{P}=\rho\mathopen{}\mathclose{{\left(E_{\rho^{-1}(P)}}}\right)\subset S$ to be the unique curve in $|{-K_{S}}|$ through $P$ and $B_{P}:=\theta(D_{P})\subset W$.
Remark 3.5.
For $P\in S^{o}$, we have the following properties.
  1. (1)
    $C_{P}\sim-3K_{S}$ [20, Lem. 3.9].
  2. (2)
    When irreducible, $C_{P}$ has geometric genus at most $1$: for general $P$, the curve $C_{P}$ has (at least) three double points at $\theta^{-1}(\theta(P))$, while $C_{P}\in|{-3K_{S}}|$ has arithmetic genus 4 by adjunction. For non-generic $P$, the geometric genus is then still bounded by $1$ by semi-continuity of the genus in flat families.
  3. (3)
    Viewing $C_{P}$ as a divisor on $S$, its restriction $C_{P}|_{D_{P}}$ to $D_{P}$ is $(2)P+Q$, where Q=[2]PQ=[-2]P on $D_{P}$ as an elliptic curve with origin $\mathcal{O}$. Indeed, $P$ appears with multiplicity at least $2$ as it is a singularity of $C_{P}$, while the expression for $Q$ follows by noting the rational equivalences $C_{P}|_{D_{P}}\sim(-3K_{S})|_{D_{P}}\sim 3\mathcal{O}$.
Proposition 3.6.
For each point $P=[x_{0}:y_{0}:z_{0}:w_{0}]\in S$ with $w_{0}\neq 0$, the set
$\{R\in S:|C_{R}\cap D_{P}|<3\}$
is a proper closed subset of $S$.
Proof.
The set is clearly closed, so it suffices to find one point $R\in S$ not lying inside it.
The curve $D_{P}$ is irreducible, thus so is $B_{P}$, which coincides with the plane section of $W\subset\mathbb{P}^{3}$ given by $\Pi_{P}:w_{0}^{3}X_{2}-w_{0}^{3}f(z_{0}/w_{0})X_{3}=0$. Let $L\subset\Pi_{P}$ be a general line. This intersects $B_{P}$ transversely in 3 points. Let $\sigma:W^{\prime}\rightarrow W$ be the composition of the minimal desingularization of $W$ and the blowup in the inverse image of $L\cap W$. Let $g:W^{\prime}\rightarrow\mathbb{P}^{1}$ be the genus-$1$ fibration corresponding to the pencil of plane sections cut out in $W$ by planes containing $L$.
Any relatively minimal model of $W^{\prime}$ is a rational elliptic surface, which has (étale/topological) Euler number $12$ [23, Ch. 5], making the Euler number $e$ of $W^{\prime}$ at least $12$. The Euler number of a regular genus-$1$ fibration is the sum $\sum_{v}e(F_{v})$ of the Euler numbers of its singular fibers $F_{v}$ (loc. cit.). Consider now the planes $\Pi_{\pm}$ passing through $L$ and $0:\pm\sqrt{c}:1:0]$, with the convention that $\Pi_{+}=\Pi_{-}$ in the case $c=0$. Let $F_{P},F_{\pm}$ be the fibers corresponding to the plane sections $\Pi_{P}\cap W$ and $\Pi_{\pm}\cap W$, respectively.
The diagram below summarizes part of what we defined so far.
[…]
We know that $B_{P}$ is irreducible. We claim that each $\Pi_{\pm}\cap W$ is also irreducible. In fact, since $L$ is a general line on $\Pi_{P}$, it is skew with respect to any finite set of lines in $\mathbb{P}^{3}$ not lying on $\Pi_{P}$. There are only finitely many lines on $W$ since it is not a cone (Lemma 3.3), and none of these lies on $\Pi_{P}$ since $\Pi_{P}\cap W=B_{P}$ is irreducible, hence $L$ is skew with respect to all lines of $W$. Then $\Pi_{\pm}\cap W$ are plane cubics with no lines, proving the claim.
From irreducibility above, we infer that $F_{P}\cong\Pi_{P}\cap W$ has $1$ component, while each $F_{\pm}$ has at most $3$ components when $F_{+}\neq F_{-}$ (equiv. $c\neq 0$) and $6$ or $7$ components when $F_{+}=F_{-}$ (equiv. $c=0$): the proper transform of $\Pi_{\pm}\cap W$ and the exceptional components coming from the desingularization (the numbers come from Lemma 3.3: $A_{k}$ and $E_{k}$-singularities give $k$ exceptional components). In each case we have $\sum_{F_{v}\in\{F_{+},F_{-}\}}e(F_{v})\leq 8$, thus $\sum_{F_{v}\in\{F_{P},F_{+},F_{-}\}}e(F_{v})\leq 10$, and there is at least one other singular fiber $F_{Q}\neq F_{P},F_{\pm}$.
Now $\sigma(F_{Q})$ is a plane section of $W$ with some singular point $R^{\prime}\neq[0:\pm\sqrt{c}:1:0]$, hence $\sigma(F_{Q})=T_{R^{\prime}}W\cap W$. Then, for $R\in\theta^{-1}(R^{\prime})$, the curve $C_{R}=\theta^{-1}(\sigma(F))\subset S$ intersects $D_{P}$ in at least 3 points. Since $C_{R}\cdot D_{P}=3$, we are done. ∎
Definition 3.7.
Let $\mathcal{C}\subset S^{o}\times S$ be the incidence variety whose $\overline{K}$-points are
$\mathcal{C}(\overline{K})=\{(P,Q)\in S^{o}(\overline{K})\times S(\overline{K}):Q\in C_{P}\}.$
Denote by $\pi_{1}:\mathcal{C}\rightarrow S^{o},\pi_{2}:\mathcal{C}\rightarrow S$ the two projections. Define the normalization morphism $r:\mathcal{C}^{\text{n}}\to\mathcal{C}$, and let $\pi^{\text{n}}_{i}:=\pi_{i}\circ r$ for $i=1,2$ be the composition.
Proposition 3.8.
The projection $\pi_{1}:\mathcal{C}\rightarrow S^{o}$ is flat and $\mathcal{C}$ is geometrically integral.
Proof.
The fibers of $\pi_{1}$ are curves in $|{-3K_{S}}|$, and in particular they all have the same Hilbert polynomial (with respect to any fixed ample line bundle on $S$), making $\pi_{1}$ flat [14, Thm. III.9.9]. For general $P$, the curve $C_{P}$ is geometrically integral [20, Prop. 3.16]. The morphism $\pi_{1}:\mathcal{C}\to S^{o}$ is now flat over a geometrically integral base with geometrically integral generic fiber, hence $\mathcal{C}$ is geometrically integral [17, Prop. 4.3.8]. ∎
Recall the blow-up map $\rho\colon\mathcal{E}\rightarrow S$.
Proposition 3.9.
The set of rational points $\mathcal{C}(K)$ on $\mathcal{C}$ is dense.
Proof.
The proof of [20, Thm. 3.1] shows that, for infinitely many $R\in\mathbb{P}^{1}(K)$, the surface $\mathcal{C}_{R}:=\mathcal{C}\times_{\pi_{1},S}D_{R}$ (i.e. the restriction of $\mathcal{C}$ to $P\in D_{R}$) has dense $K$-rational points. In particular, $\{R\in\mathbb{P}^{1}(K):(\pi\circ\rho^{-1}\circ\pi_{1})^{-1}(R)(K)\text{ is dense in }(\pi\circ\rho^{-1}\circ\pi_{1})^{-1}(R)\}$ is dense in $\mathbb{P}^{1}(K)$, thus $\mathcal{C}(K)$ is dense in $\mathcal{C}$. ∎
Proposition 3.10.
The morphism $\pi_{2}:\mathcal{C}\rightarrow S$ has geometrically integral generic fiber.
Proof.
Assume that $\pi_{2}$ does not have a geometrically integral generic fiber; then the same holds for $\pi^{\text{n}}_{2}$, and we get a relative normalization decomposition:
[…]
where $S^{\prime}$ is integral, $g$ has geometrically integral fibers, and $h$ is finite of degree at least $2$. We are going to derive a contradiction by showing that the function field extension $K(S^{\prime})/K(S)$ is a subextension of two linearly disjoint extensions of $K(S)$.
First extension. The map $\pi_{1}$ admits a rational section $S\dashrightarrow\mathcal{C},\ P\mapsto(P,Q)$ with $Q\in D_{P}$ such that Q=[2]PQ=[-2]P on the elliptic curve $D_{P}$ with origin $\mathcal{O}$ (see Remark 3.5). For general $P$, this point $Q$ has multiplicity $1$ in the intersection $D_{P}\cap C_{P}$ and is thus smooth in $C_{P}$, hence $(P,Q)$ lies in the smooth locus of $\mathcal{C}$. In particular, this section lifts to a map $s:S\dashrightarrow\mathcal{C}^{\text{n}}$. The composition $\pi^{\text{n}}_{2}\circ s:S\dashrightarrow S$ is the map $2]:S\dashrightarrow S$, and thus $K(S^{\prime})$ is contained in the degree-$4$ extension $K(S)_{2}/K(S)$ (with $K(S)_{2}\cong K(S)$ as abstract fields) given by multiplication by [2][-2].
Second extension. For general $P$, the restriction $\mathcal{C}^{n}_{P}:=\mathcal{C}^{n}\times_{\mathcal{C}}C_{P}\to C_{P}$ of the normalization morphism $\mathcal{C}^{n}\to\mathcal{C}$ above $C_{P}\cong\{P\}\times C_{P}\subset\mathcal{C}$ is the normalization morphism of $C_{P}$ and $C_{P}$ has a simple node at $P$, thus the inverse image $V$ in $\mathcal{C}^{n}$ of a sufficiently small Zariski-open $U$ of the diagonal $\Delta:=\{(P,P):P\in S^{o}\}\subset\mathcal{C}$ is an étale double cover $V\to U$. The composition
$V\to\mathcal{C}^{\text{n}}\xrightarrow{\pi^{\text{n}}_{2}}S$
is a generically finite map of degree $2$, and by (1) the field extension $K(S^{\prime})/K(S)$ of degree $\geq 2$ embeds in the étale extension $K(V)/K(S)$ of degree $2$. Hence $K(S^{\prime})=K(V)$.
Linear disjointness. Since the Mordell–Weil group of $\mathcal{E}$ is torsion-free [23, Thm. 7.4], the degree-$4$ extension $K(S)_{2}/K(S)$ given by multiplication by [2][-2] has no non-trivial proper subextensions, as any of these would have degree $2$ and would give rise to a cyclic 2-isogeny on the elliptic surface $\mathcal{E}$, hence to a non-trivial $2$-torsion section. This immediately implies linear disjointness and concludes the proof. ∎
We now prove that the surface $S$ satisfies the Hilbert property.
Proof of Theorem 1.2.
Consider the low-genus family $\pi^{\text{n}}_{2}:\mathcal{C}^{\text{n}}\to S$ with underlying fibration $\pi^{\text{n}}_{1}\colon\mathcal{C}^{\text{n}}\rightarrow S^{o}$ (this is a family with general fiber the normalization of some curve $C_{P}$ and general element $C_{P}$). We let $X$ be its inverse image in $\mathcal{E}$, i.e. $X=\mathcal{C}^{\text{n}}\times_{\pi_{2}^{\text{n}},\rho}\mathcal{E}$. The variety $X$ has a low-genus fibration with base $S^{o}$, namely the composition $X\to\mathcal{C}^{\text{n}}\xrightarrow{\pi_{1}^{\text{n}}}S^{o}$. We denote by $p_{2}$ the second projection $X\to\mathcal{E}$. By Proposition 3.10, $\pi^{\text{n}}_{2}$ has geometrically integral generic fiber, and thus the surjective base change $p_{2}$ of $\pi_{2}^{\text{n}}$ has this also.
Let $F$ denote the union of divisors on $\mathcal{E}$ which are vertical with respect to the elliptic fibration $\pi\colon\mathcal{E}\rightarrow\mathbb{P}^{1}$ and do not intersect $\rho^{-1}(C_{P})\subset\mathcal{E}$ for a general $P$ transversally. As the fibers of $\pi$ are all geometrically integral (they are Weierstrass curves), the irreducible divisors in $F$ are fibers of $\pi$. By Proposition 3.6, the only such fiber that possibly intersects the curve $C_{P}$ non-transversally for a general $P$ is the fiber $\pi^{-1}([1:0])$ at infinity. So $\mathcal{E}\setminus F$ contains $\pi^{-1}(\mathbb{A}^{1})\cong S\setminus Z$ for $Z$ the anticanonical divisor of $S$ given by $w=0$. This is simply connected, as follows for instance by applying a result of Coccia [3, Prop. 4.3(i)] (see also [14, Lem. 3.27] for the case $a_{6}(0)\neq 0$, or equivalently, smooth ZZ). [^1] Following a similar approach to Coccia’s, one may also prove this simple connectedness by first noting that the fundamental group of $S\setminus Z\cong\pi^{-1}(\mathbb{A}^{1})$ is abelian using e.g. [13, X.1.4]. Over an algebraically closed field of characteristic 0, a smooth variety with abelian fundamental group is simply connected if and only if its Picard group is torsion-free and it has no non-constant invertible functions. As the anticanonical class $[Z]$ is primitive in the torsion-free module $\operatorname{Pic}S$, both conditions hold for $S\setminus Z$.
The existence of a $P$ as in the statement gives Zariski–density of $S(K)$ by Nijgh’s work [20], and the Hilbert property for $S$ follows from Theorem 1.1. ∎
To conclude, we prove Theorem 1.5.
Proof of Theorem 1.5.
Let $U\subset\mathbb{A}_{K}^{9}$ be the nonempty Zariski-open whose $\bar{K}$-points are $9$-tuples $(a,b,\ldots,f_{3})$ corresponding to a smooth $S$, and let $X\subset U\times\mathbb{P}(2,3,1,1)$ be the total space associated to the parametric family of Definition 3.1. Then the projection $\pi=pr_{1}:X\to U$ is smooth surjective and its fibers are del Pezzo surfaces of degree $1$. Note that $X$ is birational to the affine hypersurface $Y$ in $\mathbb{A}_{K}^{11}$ defined by the equation
$y^{2}=x^{3}+(af(z)+b)x+(cf(z)^{2}+df(z)+e).$
The hypersurface $Y$ is rational via projection to any coordinate appearing linearly.
For $N\in\mathbb{Z}_{\geq 1}$, let $V_{N}\subset X$ be the non-empty Zariski-open corresponding to pairs $(\underline{a},P),\underline{a}=(a,\ldots,e,f_{0},\ldots,f_{3}),P=[x_{0}:y_{0}:z_{0}:w_{0}]$ such that $w_{0}\neq 0,3z_{0}+2f_{2}w_{0}/f_{3}\neq 0,$ the polynomial $f(t)-f(z_{0}/w_{0})$ is separable, and $N](\rho^{-1}(P))\neq O$ on $\mathcal{E}_{P}$. Choose now $N$ to be divisible by the torsion exponents of all elliptic curves over $K$. This is a finite number by (the generalization of) Merel’s theorem.
For any $x\in V_{N}(K)$, the del Pezzo surface $\pi^{-1}(\pi(x))$ has the Hilbert property by Theorem 1.2. The morphism $\pi|_{V_{N}}:V_{N}\to U$ has geometrically integral generic fiber and thus sends non-thin sets to non-thin sets. Then, since $V_{N}(K)$ is non-thin by Hilbert’s Irreducibility Theorem (recall that $X$, and thus $V_{N}$, is rational), $\pi(V_{N}(K))$ is non-thin.
Finally, for $\eta$ the generic point of $U$, we have the following properties for $X\to U$:
  1. (1)
    for every $u\in U$, we have $\operatorname{rk}\operatorname{Pic}X_{u}\geq\operatorname{rk}\operatorname{Pic}X_{\eta}$;
  2. (2)
    the set of $u\in U$ such that $\operatorname{rk}\operatorname{Pic}X_{u}>\operatorname{rk}\operatorname{Pic}X_{\eta}$ is thin.
The first property follows from injectivity of specialization on the Picard group (see e.g. [29, §5]). More precisely, in (2) we mean that there exists a finite collection of irreducible covers $\pi_{i}:Y_{i}\to U,i=1,\ldots,r$ of finite degrees >1>1 such that $\operatorname{rk}\operatorname{Pic}X_{u}>\operatorname{rk}\operatorname{Pic}X_{\eta}$ if and only if $u\in\operatorname{im}(Y_{i}(\kappa(u))\to U(\kappa(u)))$. This holds by applying Hilbert’s Irreducibility Theorem to the family of degree-240 polynomials whose roots parametrize the lines in each surface in our family. The Galois orbits of lines correspond to the irreducible factors of these polynomials and determine the Picard rank. The Picard rank can only increase if an irreducible factor becomes reducible upon specialization, which occurs over a thin locus. By a result of Kloosterman [16, Lem. 2.5, Prop. 2.9], reproving a result of Desjardins and Naskrecki [10, Sec. 4.2]), the del Pezzo surface
$y^{2}=x^{3}+Az^{6}+Bw^{6},$
where we take $A$ and $B$ to be transcendental variables over $K$, has Picard rank $1$. This surface is a member of the family $X\to U$: it is the specialization at the generic point $\xi$ of the copy of $\mathbb{A}^{2}_{K}$ in $\mathbb{A}^{9}_{K}$ defined by $f_{3}=A,e=B,c=1$, and where all the other coordinates are trivial. Then by (1) we get $1\leq\operatorname{rk}\operatorname{Pic}X_{\eta}\leq\operatorname{rk}\operatorname{Pic}X_{\xi}=1$, thus $\operatorname{rk}\operatorname{Pic}X_{\eta}=1$. Hence by (2) the set $T\subset U(K)$ corresponding to del Pezzo surfaces with Picard rank $\geq 2$ is thin. Then $\pi(V_{N}(K))\setminus T$ furnishes the sought non-thin set of parameters. ∎